OpenStudy (anonymous):

Simplify... b2/3a(exponent -4)

OpenStudy (anonymous):

Is this your question? \[\frac{b^2}{3a^{-4}}\]

OpenStudy (anonymous):

Or this? \[\frac{b^2}{(3a)^{-4}}\]

OpenStudy (anonymous):

Either way, the negative exponent simply means to move the term to the opposite side of the fraction. \(\Large{x^{-1} = \frac{1}{x}}\)

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Your Welcome.